# Covariant formulation of electrodynamics in isotropic media revisited

### v3

@@ 2 Some basic results of covariant electrodynamics in the vacuum
Recall that $\varphi$ and $\vec{A}$ must satisfy the simplified equations (12) but also the Lorenz gauge; the last condition is fulfilled if the continuity equation is satisfied. The potentials $\varphi$ and $\vec{A}$ that satisfy the simplified equations (12) and the Lorenz gauge, satisfy also the inhomogeneous Maxwell’s equations in the vacuum (11).
An analogous remark is also valid if an arbitrary gauge is introduced into eqs. (11): $\varphi$ and $\vec{A}$ must satisfy the corresponding simplified equations but also the gauge chosen; the last condition is fulfilled if the continuity equation is satisfied [9].
@@
For example, one can use the Coulomb–gauge four–potential $A^{\mu}_{\mbox{\tiny C}}=(\varphi_{\mbox{\tiny C}},\vec{A}_{\mbox{\tiny C}})$ in a frame $S$ and Lorentz–transform it to another frame $S^{\prime}$ to obtain $A^{\prime\mu}_{\mbox{\tiny C}}=(\varphi^{\prime}_{\mbox{\tiny C}},\vec{A}^{\prime}_{\mbox{\tiny C}})$. Since the Coulomb gauge condition $\vec{\nabla}\cdot\vec{A}_{\mbox{\tiny C}}=0$ is not covariant, $\vec{\nabla}^{\prime}\cdot\vec{A}^{\prime}_{\mbox{\tiny C}}$ does not vanish. But, $\vec{E}$ and $\vec{B}$ are gauge invariant, so one obtains the same $\vec{E}$ and $\vec{B}$ in the original $S$ frame either using the Coulomb gauge or a covariant gauge such as the Lorenz gauge, $\partial_{\mu}A^{\mu}_{\mbox{\tiny L}}=0$, see, e.g., [9]. On the other hand, since the field equations written in terms of the potentials do obey a covariant set of equations, eqs. (8) and (11), one obtains the same $\vec{E}^{\prime}$ and $\vec{B}^{\prime}$ in $S^{\prime}$, regardless of which gauge is used in the $S$ frame (the gauge employed need not be covariant). An instructive example is found in Reference [10].

### v4

@@ 2 Some basic results of covariant electrodynamics in the vacuum
Recall that $\varphi$ and $\vec{A}$ must satisfy the simplified equations (12) but also the Lorenz gauge; for the retarded potentials, the last condition is fulfilled if and only of the continuity equation is satisfied. The potentials $\varphi$ and $\vec{A}$ that satisfy the simplified equations (12) and the Lorenz gauge, satisfy also the inhomogeneous Maxwell’s equations in the vacuum (11).
An analogous remark is also valid if an arbitrary gauge is introduced into eqs. (11): $\varphi$ and $\vec{A}$ must satisfy the corresponding simplified equations but also the gauge chosen; for an explicit retarded solution of the simplified equations, the last condition is fulfilled if and only if the continuity equation is satisfied [9].
@@
For example, one can use the Coulomb–gauge four–potential $A^{\mu}_{\mbox{\tiny C}}=(\varphi_{\mbox{\tiny C}},\vec{A}_{\mbox{\tiny C}})$ in a frame $S$ and Lorentz–transform it to another frame $S^{\prime}$ to obtain $A^{\prime\mu}_{\mbox{\tiny C}}=(\varphi^{\prime}_{\mbox{\tiny C}},\vec{A}^{\prime}_{\mbox{\tiny C}})$. Since the Coulomb gauge condition $\vec{\nabla}\cdot\vec{A}_{\mbox{\tiny C}}=0$ is not covariant, $\vec{\nabla}^{\prime}\cdot\vec{A}^{\prime}_{\mbox{\tiny C}}$ does not vanish. But, $\vec{E}$ and $\vec{B}$ are gauge invariant, so one obtains the same $\vec{E}$ and $\vec{B}$ in the original $S$ frame either using the Coulomb gauge or a covariant gauge such as the Lorenz gauge, $\partial_{\mu}A^{\mu}_{\mbox{\tiny L}}=0$, see, e.g., [9]. Now since the field equations written in terms of the potentials do obey a covariant set of equations, eqs. (8) and (11), one obtains the same $\vec{E}^{\prime}$ and $\vec{B}^{\prime}$ in $S^{\prime}$, regardless of which gauge is used in the $S$ frame (the gauge employed need not be covariant). Shortly, while $A^{\mu}_{\mbox{\tiny C}}\neq A^{\mu}_{\mbox{\tiny L}}$, due to gauge invariance $F^{\mu\nu}_{\mbox{\tiny C}}=F^{\mu\nu}_{\mbox{\tiny L}}$ and thus $F^{\prime\mu\nu}_{\mbox{\tiny C}}=F^{\prime\mu\nu}_{\mbox{\tiny L}}$. An instructive example is found in Reference [10].